(3) At current factor prices and given its current level of capital, if it sets price equal to short run marginal cost it chooses QR- 20.ISMC(SR,w.r,K) 10, SAVC(Q, w,r, K) 8, and SAFC(Q, w,r,K) -9, what is the firms profits in the short run? Will it shut down or operate in the short run?
Answer #1

Q= 20

SMC=10

SAFC= 9

SAVC=8

SATC= SAVC+SAFC= 9+8=17

Loss = (SATC-P)*Q = (17-10)*20= 140

loss =140 or profit = -140.

Since P=MC< SRTC but P>SAVC the firm will operate in the short run.